# Tokenizing Buildings: A Transformer for Layout Synthesis

### v1

@@ 4 Experiments and Results
We structure this section around our two main parts, DDEP and SBM embeddings, and compare our results to previous approaches. Due to space constraints, we defer training, implementation details and full tokenizer/adapter configurations to the Supplementary Material. In all experiments, we use DDEP in its unguided form: the decoder autoregressively generates entity sequences conditioned only on the room envelope and encoder memory, without retrieval-based conditioning or external agent feedback.
@@
Coverage Navigability OC (%) Inference Latency (s) Method $\uparrow$ mean $\pm$ std $\uparrow$ mean $\pm$ std $\uparrow$ SR (mean $\pm$ std) $\downarrow$ DF (mean $\pm$ std) $\downarrow$ mean $\pm$ std $\downarrow$ mean $\pm$ std Text-based methods Claude Sonnet 4.5 34.1 ± 0.3 2.1 ± 2.7 0.26 ± 0.02 0.69 ± 0.03 33.3 ± 2.5 18.02 ± 12.19 Gemini 2.5 Pro 67.1 ± 1.6 58.0 ± 8.3 0.69 ± 0.07 0.30 ± 0.03 10.2 ± 1.6 42.51 ± 13.25 GPT-5 Medium 56.8 ± 1.5 51.2 ± 8.8 0.63 ± 0.07 0.32 ± 0.04 12.9 ± 1.8 100.86 ± 65.58 Qwen3 60.5 ± 1.4 34.0 ± 0.8 0.50 ± 0.01 0.45 ± 0.02 18.1 ± 1.6 21.27 ± 13.32 VLM methods Claude Sonnet 4.5 44.9 ± 3.4 52.7 ± 1.7 0.64 ± 0.01 0.32 ± 0.03 23.7 ± 2.0 18.71 ± 3.62 Gemini 2.5 Pro 73.2 ± 1.2 54.6 ± 2.2 0.67 ± 0.02 0.34 ± 0.02 11.9 ± 1.9 43.16 ± 13.16 GPT-5 Medium 55.5 ± 1.4 54.3 ± 10.4 0.65 ± 0.08 0.29 ± 0.05 12.3 ± 3.9 164.43 ± 105.88 GPT-5 High 58.3 ± 0.0 57.8 ± 54.0 0.67 ± 0.42 0.27 ± 0.39 12.1 ± 25.5 124.18 ± 78.34 Qwen3 75.9 ± 1.5 48.7 ± 1.0 0.60 ± 0.02 0.33 ± 0.02 12.4 ± 1.6 439.04 ± 377.43 Domain-specific methods LayoutVLM 40.3 ± 59.8 0.55 ± 0.45 0.41 ± 0.45 4.9 ± 11.4 144.99 ± 36.50 FlairGPT 46.6 ± 36.5 28.2 ± 61.0 0.46 ± 0.46 0.51 ± 0.46 5.0 ± 12.9 1133.61 ± 1596.40 DDEP (Ours) 98.2 ± 6.3 82.4 ± 40.7 0.89 ± 0.31 0.18 ± 0.29 3.2 ± 10.2 3.18 ± 0.18
Table 1: Comparison on the held-out set. We report mean $\pm$ std. Higher is better for Coverage, Navigability (mean and SR); lower is better for OC and DF. Best overall (SOTA) also in bold. LayoutVLM requires an initial furniture list as input, so it is excluded from Coverage metric. Latency is reported as mean $\pm$ std per room, measured end-to-end on an AWS EC2 g5.xlarge instance.
4.1 Experimental Setup
4.1.1 Dataset and Setup
We train SBM on a proprietary corpus of residential BIM scenes containing single-family homes with typed rooms (e.g., bathrooms, dining rooms, offices, pantries). Across these constraints, the dataset contains 439,932 training samples over. We construct held-out splits per room type and report all quantitative results on the same test rooms for every method.
Each room is converted to BIM-Token Bundles and then to encoder and decoder sequences as in Sec. 3.2, yielding relatively short sequences compared to text corpora. Over all room types and training regimes, SBM sees on the order of five million tokens in total across encoder and decoder streams, i.e., a few million envelope tokens and a comparable number of entity tokens. Despite this modest scale, the dataset covers a wide range of residential geometries and furniture programs, providing enough diversity to train both the encoder-only embedding pathway and the DDEP decoder for the room types evaluated in Secs. 4.2 and 4.4.
4.1.2 Compared Methods
We evaluate three families of methods on the same held-out rooms. For each room envelope, we construct a serialized text description and, when applicable, a rendered floorplan image. All methods return predicted props and casework, which we map to a unified schema of entity categories and dimensions. For LLM/VLM baselines we use fixed prompts (see Suppl.) and decoding (T=0.7, top‑p=0.9 unless stated). VLMs condition on a 1024×1024 floorplan render (orthographic, 20px/m). Free‑text outputs are parsed via a rule‑based mapper to our inventory. Each method runs 3 seeds; we report mean +/- sd and inference time. For reproducible implementation details, refer to Supplementary Material.
Text LLMs. Claude Sonnet 4.5, Gemini 2.5 Pro, GPT-5, and Qwen3 in text-only mode.
Vision–language models (VLMs). The same base models with image-conditioned variants that receive a floorplan render plus textual context.
Domain-specific methods. LayoutVLM [sun2025layoutvlm], FlairGPT [littlefair2025flairgpt], and our DDEP model built on the SBM encoder–decoder backbone.
4.1.3 Layout Quality Metrics
We report three complementary layout metrics: coverage (inventory satisfaction), navigability (functional access), and overlap & clearance (spatial violations). All metrics use the same room geometry, entity dimensions, and inventory specification for every method. Full implementation details are provided in the Supplement.
Coverage.
Coverage measures how well a predicted layout $\hat{y}$ satisfies the room-type inventory $y$, independent of exact placement. Let $\mathcal{I}$ be required items and $\mathcal{G}$ mutually exclusive alternative groups with non-negative weights $w_{i},w_{g}$:
$\mathrm{Cov}(y,\hat{y})=\frac{\sum_{i\in\mathcal{I}}w_{i}\,s_{i}(\hat{y})\;+\;\sum_{g\in\mathcal{G}}w_{g}\,s_{g}(\hat{y})\;-\;p_{\text{extra}}(\hat{y})}{\sum_{i\in\mathcal{I}}w_{i}\;+\;\sum_{g\in\mathcal{G}}w_{g}}.$
Item scores $s_{i}\!\in[0,1]$ handle missing/overfill counts; group scores $s_{g}\!\in[0,1]$ capture satisfaction of any valid alternative; $p_{\text{extra}}\!\geq\!0$ penalizes extraneous categories.
Navigability.
The navigability metric evaluates the walkability from door-to-target. Let $\mathcal{P}$ denote all door–target pairs, and let $\ell_{t}$ be the shortest collision-free path length for pair $t$ (computed with A*), with $\ell_{t}^{\star}$ the straight-line distance inside the room. We define the success rate and detour factor as $\mathrm{SR}=\frac{1}{|\mathcal{P}|}\sum_{t\in\mathcal{P}}\mathbf{1}[\ell_{t}<\infty],\quad\mathrm{DF}=\frac{1}{|\mathcal{P}_{\text{reach}}|}\sum_{t\in\mathcal{P}_{\text{reach}}}\frac{\ell_{t}}{\ell_{t}^{\star}},$ where $\mathcal{P}_{\text{reach}}=\{t\in\mathcal{P}:\ell_{t}<\infty\}$. For visualization we also report a scalar navigability score $\mathrm{Nav}=100\times\big(\mathrm{SR}-\lambda\,\mathrm{DF}\big)$, with $\lambda$ chosen on a validation set to balance reachability and detour.
Overlap & clearance.
The overlap & clearance metric measures spatial violations using exact polygon geometry. We compute four normalized terms: entity overlap fraction (EOF), global overlap area (GOA), door clearance intrusion (DCI), and wall-bounds violation (WBV). These are combined into a single overlap/clearance score
$\mathrm{OC}=w_{\text{EOF}}\,\mathrm{EOF}+w_{\text{GOA}}\,\mathrm{GOA}+w_{\text{DCI}}\,\mathrm{DCI}+w_{\text{WBV}}\,\mathrm{WBV},$
with fixed weights $w_{\cdot}$ shared for all methods.
4.2 DDEP Quantitative Comparisons
@@
On the held-out benchmark (Table 1), we compare layout generators using three metrics. Text LLMs achieve moderate coverage but weak functional quality (under-furnishing, low reachability, higher OC). VLMs improve some axes but exhibit clear trade-offs (gains in coverage often accompany reduced reachability or more collisions). Domain-specific baselines (LayoutVLM, FlairGPT) show lower violation rates yet remain weak on both coverage and navigability under our unified schema. DDEP occupies a distinct regime: near-complete coverage, the highest navigability, and the lowest OC—despite placing substantially more entities than any baseline. No competing method attains high coverage and high navigability simultaneously without incurring larger spatial errors, positioning DDEP as state of the art on this benchmark.
@@ 4.3 Qualitative Comparisons
Figure 3 shows example layouts for several room types from DDEP, VLMs, and domain-specific methods. VLM baselines often overfill rooms with extra casework and furniture, blocking circulation and door clearances, especially in living rooms and garages. Domain-specific methods tend to place more appropriate furniture types but still leave narrow or blocked paths around key elements. In contrast, DDEP produces layouts that satisfy the room program while keeping clear, door-connected circulation bands and respecting wall-referenced anchoring, visually matching the trends reported in our quantitative metrics.
@@ 4.4 Embedding Evaluation and Encoder Ablations
We evaluate SBM’s encoder-only pathway for room embedding quality through retrieval and clustering benchmarks. As an ablation study, we compare against state-of-the-art text embedding baselines to demonstrate SBM’s encoder capacity and highlight its complementary capabilities for geometric and spatial understanding. We compare against E5-Large-v2 [wang2022e5]: 560M parameters, pre-trained on Common Crawl and Wikipedia, BGE-Large-en-v1.5 [bge_embedding]: 335M parameters, pre-trained on web-scale text data, and GTE-Large-en-v1.5 [li2023towards]: 335M parameters, pre-trained on web-scale text corpora. These baselines are pre-trained on billions of text tokens, representing orders of magnitude more training data than SBM’s domain-specific dataset. For fair comparison, we serialize room data to text format, enabling text embeddings to process the same room information.
@@
We assess retrieval quality with nDCG@k (graded relevance) and Success@k/MRR/Precision@k, structural consistency with type-constrained nDCG@k and triplet accuracy, entity-level similarity via overlap of props/casework/doors, and global organization with clustering metrics (NMI/ARI/Silhouette) and geometric rank correlation (Spearman/Kendall) against a geometry oracle. Full metric definitions and cross-type retrieval tables are provided in the Supplement; here we focus on type-constrained retrieval and clustering/correlation.
Method nDCG Triplet Entity Sim Clust.(NMI) SBM 62.6 100.0 78.9 0.640 E5-Large-v2 86.7 99.9 78.4 0.371 BGE-Large-en-v1.5 80.1 99.9 77.0 0.358 GTE-Large-en-v1.5 74.3 99.8 77.0 0.335
Table 2: Type-Constrained Retrieval Performance @ k=5
Method NMI ARI Silhouette Spearman SBM 0.640 0.307 0.152 0.145 E5-Large-v2 0.371 -0.019 0.051 0.315 BGE-Large-en-v1.5 0.358 -0.003 0.069 0.235 GTE-Large-en-v1.5 0.335 -0.002 0.067 0.133
Table 3: Clustering and Geometric Correlation Metrics
Table 2 shows that text encoders lead on ranking quality within type (nDCG: E5 86.7, BGE 80.1, GTE 74.3 vs. SBM 62.6), while SBM excels at structure: it attains perfect triplet accuracy (100.0), the highest entity overlap (78.9), and stronger type organization (NMI 0.640). In short, text models rank better among same-type rooms, whereas SBM preserves within-type geometry and entities more faithfully.
Figure 4: UMAP visualization of room embeddings colored by room type category. SBM embeddings (left, NMI: 0.640) exhibit well-separated clusters with distinct boundaries between room types, demonstrating superior geometric and spatial understanding. E5-Large-v2 embeddings (right, NMI: 0.371) show more intermingled clusters with blurred boundaries, indicating weaker room type separation. The 1.7× higher NMI score for LBM reflects its specialization in capturing geometric structure and spatial relationships inherent in building layouts, rather than semantic similarity alone.
As summarized in Table 3, SBM yields markedly better global organization (NMI/ARI/Silhouette: 0.640/0.307/0.152) than text baselines (NMI 0.335–0.371), but aligns less with our geometry oracle in rank correlation (Spearman 0.145 vs. E5’s 0.315). This indicates a tighter type-structured space for SBM (Fig. 4) but stronger oracle-ranked agreement for E5.

### v2

@@ 4 Experiments and Results
We evaluate DDEP and SBM embeddings on a proprietary corpus of residential BIM scenes (single-family homes with typed rooms). Training, implementation details, and full tokenizer/adapter configurations are in the Supplementary Material. We report DDEP results under two complementary protocols whose rows are not directly comparable.
Frontier LLM Benchmark. Production DDEP is trained per room type on the full augmented corpus (439,932 samples) and compared to frontier LLMs, VLMs, and domain-specific baselines under deployment-realistic conditions on a 50-room held-out set.
Controlled Same-Data Benchmark. A frozen split of 15,740/2,009/1,826 (train/val/test) rooms across five types (bathfull, bedroom, laundry, living, master_bed) with shared ontology normalization, adapters, and evaluation harness for all methods. The camera-ready metric pipeline first filters this test split to the 1,231-room geometry-valid subset, and Table 2 reports the shared 1,225-room intersection on which all compared methods materialize outputs under this harness. This protocol isolates architectural differences under matched data conditions.
Metrics.
We report three complementary metrics, all computed with identical room geometry, entity dimensions, and inventory specifications across methods. Coverage ($\uparrow$) measures inventory satisfaction: the weighted fraction of required items and alternative groups fulfilled, penalizing extraneous categories. Navigability ($\uparrow$) evaluates functional access via A*-based pathfinding from inward-offset door portals to essential targets. Doorways must remain valid entry points: if the portal or immediate inward vestibule is blocked, that door contributes failed door-target pairs except for a tiny raster snap used only for numerical robustness. For reachable pairs we compare shortest-path length $\ell_{t}$ to straight-line distance $\ell_{t}^{\star}$ through a capped detour penalty, and report $\mathrm{Nav}=100\times(\mathrm{SR}-\lambda\,\mathrm{DF})$ with $\lambda=0.35$. When nothing is reachable, $\mathrm{DF}$ defaults to its worst value and Nav reaches its minimum of 35-35. Overlap & Clearance (OC, $\downarrow$) reports the weighted composite of entity overlap fraction, global overlap area, and door-clearance intrusion; wall-bounds violation is tracked separately in the Supplement. Full definitions, including Coverage formulation and OC weights, are in the Supplement.
Table 1: Frontier LLM Benchmark. Comparison of frontier LLMs on the 50-room held-out set. We report mean $\pm$ std. Higher is better for Coverage and Navigability (mean and SR); lower is better for OC, DF, and latency. OC reports the weighted overlap-clearance composite defined in the Supplement. Best text/VLM method per metric in bold among those evaluated on all 50 rooms; subset-evaluated models are excluded from bolding. All methods use the same unguided prompt with entity catalog (information parity.) Evaluated on a subset of rooms due to low parse success rate (<<70%). Evaluated on 37/50 rooms. §Evaluated on 7/50 rooms (unreliable). LayoutVLM requires an initial furniture list as input, so it is excluded from Coverage.
@@
Coverage Navigability OC (%) Inference Latency (s) Method $\uparrow$ mean $\pm$ std $\uparrow$ mean $\pm$ std $\uparrow$ SR (mean $\pm$ std) $\downarrow$ DF (mean $\pm$ std) $\downarrow$ mean $\pm$ std $\downarrow$ mean $\pm$ std Text-based methods Claude Opus 4.6 58.9 $\pm$ 37.0 -0.8 $\pm$ 56.3 0.26 $\pm$ 0.43 0.75 $\pm$ 0.40 11.3 $\pm$ 11.1 15.6 GPT-5.2 47.6 $\pm$ 36.3 16.8 $\pm$ 54.7 0.35 $\pm$ 0.40 0.53 $\pm$ 0.47 13.0 $\pm$ 13.5 55.1 Claude Sonnet 4.6 64.9 $\pm$ 35.9 0.3 $\pm$ 53.3 0.25 $\pm$ 0.39 0.70 $\pm$ 0.43 10.5 $\pm$ 11.5 81.2 Gemini 3.1 Pro 63.7 $\pm$ 37.2 20.5 $\pm$ 60.7 0.40 $\pm$ 0.46 0.56 $\pm$ 0.46 9.2 $\pm$ 14.0 67.1 Codex 5.3 68.3 $\pm$ 34.2 18.9 $\pm$ 55.8 0.37 $\pm$ 0.42 0.52 $\pm$ 0.46 9.8 $\pm$ 11.4 20.8 Claude Haiku 4.5 55.3 $\pm$ 38.3 4.4 $\pm$ 56.7 0.28 $\pm$ 0.43 0.69 $\pm$ 0.43 15.2 $\pm$ 14.5 7.0 Qwen 3.5 64.6 $\pm$ 31.5 3.1 $\pm$ 51.7 0.26 $\pm$ 0.37 0.66 $\pm$ 0.44 13.6 $\pm$ 15.5 19.0 VLM methods Claude Opus 4.6 61.9 $\pm$ 36.7 55.5 $\pm$ 56.1 0.66 $\pm$ 0.44 0.29 $\pm$ 0.40 8.2 $\pm$ 11.7 15.3 GPT-5.2 57.0 $\pm$ 38.3 34.9 $\pm$ 60.3 0.51 $\pm$ 0.46 0.46 $\pm$ 0.47 12.1 $\pm$ 11.8 101.7 Claude Sonnet 4.6 65.6 $\pm$ 35.3 34.6 $\pm$ 58.7 0.50 $\pm$ 0.45 0.43 $\pm$ 0.44 11.0 $\pm$ 13.2 14.8 Gemini 3.1 Pro§ 86.5 $\pm$ 12.6 47.8 $\pm$ 65.3 0.61 $\pm$ 0.49 0.38 $\pm$ 0.48 2.2 $\pm$ 3.0 Claude Haiku 4.5 60.7 $\pm$ 36.1 25.2 $\pm$ 59.6 0.42 $\pm$ 0.46 0.49 $\pm$ 0.46 16.2 $\pm$ 15.7 6.9 Qwen 3.5 60.7 $\pm$ 36.5 27.1 $\pm$ 54.8 0.42 $\pm$ 0.41 0.42 $\pm$ 0.46 12.8 $\pm$ 15.9 84.1 Domain-specific methods LayoutVLM 40.3 $\pm$ 59.8 0.55 $\pm$ 0.45 0.41 $\pm$ 0.45 3.8 $\pm$ 7.3 144.99 $\pm$ 36.50 FlairGPT 46.6 $\pm$ 36.5 28.2 $\pm$ 61.0 0.46 $\pm$ 0.46 0.51 $\pm$ 0.46 3.9 $\pm$ 8.1 1133.61 $\pm$ 1596.40 DDEP (Ours) 98.1 $\pm$ 6.3 79.2 $\pm$ 44.7 0.86 $\pm$ 0.34 0.20 $\pm$ 0.32 1.9 $\pm$ 3.1 3.18 $\pm$ 0.18
Table 2: Controlled Same-Data Benchmark. Controlled benchmark on the frozen same-data split, reported on the shared 1,225-room intersection of successfully materialized outputs across all methods. All methods use canonical furniture dimensions and a shared denormalization pipeline with unified metrics computation. OC reports the weighted overlap-clearance composite defined in the Supplement. Inference latency is omitted in this controlled benchmark. Rows in this table are not directly comparable to Table 1 due to different training protocols. Nav can be negative because unreachable rooms receive the worst detour penalty under the doorway-validity definition (see Sec. 4).
Coverage Navigability OC (%) Method $\uparrow$ mean $\pm$ std $\uparrow$ mean $\pm$ std $\uparrow$ SR (mean $\pm$ std) $\downarrow$ DF (mean $\pm$ std) $\downarrow$ mean $\pm$ std ATISS [Paschalidou2021NEURIPS] (same-data) 40.8 $\pm$ 23.6 -30.3 $\pm$ 24.0 0.03 $\pm$ 0.18 0.97 $\pm$ 0.18 0.5 $\pm$ 2.5 BLT [10.1007/978-3-031-19790-1_29] (same-data) 17.8 $\pm$ 20.4 -18.6 $\pm$ 42.7 0.12 $\pm$ 0.32 0.88 $\pm$ 0.31 35.7 $\pm$ 16.9 DDEP (Ours) 54.9 $\pm$ 38.3 14.6 $\pm$ 56.0 0.36 $\pm$ 0.42 0.60 $\pm$ 0.44 13.4 $\pm$ 12.5 DDEP ablations    Wall-Only 50.8 $\pm$ 39.6 9.6 $\pm$ 54.8 0.32 $\pm$ 0.41 0.63 $\pm$ 0.44 12.7 $\pm$ 11.8    No Continuous Groups 42.6 $\pm$ 40.2 0.7 $\pm$ 52.2 0.25 $\pm$ 0.39 0.69 $\pm$ 0.43 13.4 $\pm$ 13.8    No Edge Coordinates 56.7 $\pm$ 39.7 14.4 $\pm$ 57.6 0.35 $\pm$ 0.43 0.60 $\pm$ 0.45 14.5 $\pm$ 12.7    Order: t_value First 42.6 $\pm$ 39.1 4.1 $\pm$ 54.5 0.29 $\pm$ 0.41 0.70 $\pm$ 0.41 10.6 $\pm$ 12.0    Order: edge,t_value Desc 46.5 $\pm$ 39.9 7.8 $\pm$ 54.8 0.31 $\pm$ 0.41 0.66 $\pm$ 0.42 13.4 $\pm$ 13.2
4.1 Frontier LLM Benchmark
@@
We compare production DDEP against frontier LLMs and VLMs (text-only and image-conditioned variants of Claude Opus/Sonnet 4.6, Claude Haiku 4.5, GPT-5.2, Gemini 3.1 Pro, Codex 5.3, and Qwen 3.5) and domain-specific baselines (LayoutVLM [sun2025layoutvlm], FlairGPT [littlefair2025flairgpt]), all used off-the-shelf (inference only) with our room representation adapted to each method’s input format. All baselines receive the same unguided prompt with entity catalog; VLMs additionally condition on a $\times1024$ floorplan render. Free-text outputs are parsed via a rule-based mapper to our inventory (details in Suppl.). Because frontier API costs limit the test set to 50 rooms, this benchmark characterizes deployment-realistic performance across many baselines; the controlled same-data benchmark (Sec. 4.2) provides the larger-scale statistical analysis.
Table˜1 shows that production DDEP achieves 98.1 coverage, 79.2 navigability, and 1.9% OC, outperforming all baselines on coverage, navigability, and composite overlap-clearance simultaneously. Under the stricter doorway-validity metric, frontier performance drops substantially: the strongest fully evaluated text model is Codex 5.3 with 68.3 coverage and 18.9 navigability, while the strongest fully evaluated VLM is Claude Opus 4.6 with 55.5 navigability and Claude Sonnet 4.6 with 65.6 coverage. Domain-specific baselines (FlairGPT, LayoutVLM) narrow the OC gap (3.9% and 3.8%, respectively) but remain substantially weaker on coverage and navigability, indicating that DDEP’s BIM-native tokenization and wall-referenced decoding jointly enable the strongest balance across all three axes.
4.2 Controlled Same-Data Benchmark
We compare DDEP against ATISS [Paschalidou2021NEURIPS], BLT [10.1007/978-3-031-19790-1_29], and five DDEP ablations—eight models trained from scratch on the frozen controlled split under a uniform, resource-matched protocol: a fixed budget of 200 training epochs, shared ontology normalization, canonical furniture dimensions, and a common evaluation harness with no method-specific tuning.
Baselines.
DDEP achieves 54.9 coverage and 14.6 navigability on the shared 1,225-room controlled subset (Table˜2), outperforming ATISS by +14.1 coverage and +44.9 navigability and BLT by +37.1 coverage and +33.2 navigability. ATISS was designed for 3D-FRONT–style scenes with axis-aligned bounding boxes; we decode its outputs with its published descale formula, affine-map positions into the physical room, and project to the nearest wall edge to obtain wall-referenced coordinates (full pipeline in Suppl.). BLT targets 2D graphic-design layout; its $(x,y,w,h)$ outputs are similarly projected. Both baselines use their published default hyperparameters with no method-specific tuning. The resulting bridge is shared between baselines, so it cannot explain DDEP’s advantage. ATISS’s near-zero OC (0.5%) should not be read as superior spatial validity: many ATISS outputs are sparse or nearly empty, which trivially avoids collisions but yields strongly negative navigability. BLT shows the complementary failure mode, with 35.7% OC and only 0.12 SR. DDEP is the only method that maintains nontrivial coverage, positive navigability, and moderate OC under the same-data protocol.
[…]
Ablation analysis.
Five ablations probe key design choices (Table˜2, bottom). We report deltas relative to the full DDEP model (54.9 coverage, 14.6 navigability). The ordering is stable across the shared 1,225-room subset: removing continuous-group embeddings is the most damaging change, while removing explicit edge coordinates is nearly neutral on coverage and navigability but still hurts overlap-clearance.
No Continuous Groups (-12.3 cov, -13.9 nav) replaces the multi-continuous embedding which jointly embeds correlated continuous features (e.g., edge coordinates, thickness pairs) via self-attention—with independent scalar MLPs. This is the most damaging single change, confirming that joint embedding of correlated geometric attributes is critical for spatial reasoning.
Ordering ablations. t_value First (-12.3 cov, -10.5 nav) reorders entities by wall-referenced position before edge index; edge,t_value Desc (-8.4 cov, -6.8 nav) reverses the within-wall sort direction. Both confirm that the canonical edge-then-position ordering provides an inductive bias the model relies on, with edge grouping mattering most.
Wall-Only (-4.1 cov, -5.0 nav) strips doors and windows from the encoder. The moderate drop confirms that opening positions help the decoder reason about clearances, though walls alone provide a useful baseline signal.
No Edge Coordinates (+1.8 cov, -0.2 nav) removes wall endpoint coordinates while retaining lengths, thicknesses, and opening parameters. The near-tie suggests the model infers relative geometry from remaining features; however, OC rises from 13.4% to 14.5%. Per-room-type breakdowns are in the Supplement.
Notably, the two largest ablation effects—joint continuous embedding and entity ordering—each cause double-digit coverage drops, indicating that these representation design choices, not model scale or architecture depth alone, are primary drivers of DDEP’s advantage.
@@ 4.3 Qualitative Comparisons
Figures˜1 and 3 shows example layouts for several room types from DDEP, VLMs, and domain-specific methods. VLM baselines often overfill rooms with extra casework and furniture, blocking circulation and door clearances, especially in living rooms and garages. Domain-specific methods tend to place more appropriate furniture types but still leave narrow or blocked paths around key elements. In contrast, DDEP produces layouts that satisfy the room program while keeping clear, door-connected circulation bands and respecting wall-referenced anchoring, visually matching the trends reported in our quantitative metrics.
@@ 4.4 Embedding Evaluation and Encoder Ablations
We evaluate SBM’s encoder-only pathway for room embedding quality through retrieval and clustering benchmarks, comparing against three text embedding baselines: E5-Large-v2 [wang2022e5] (560M params), BGE-Large-en-v1.5 [bge_embedding] (335M), and GTE-Large-en-v1.5 [li2023towards] (335M). All three are pre-trained on web-scale corpora (orders of magnitude more data than SBM’s domain-specific dataset). For fair comparison, we serialize room data to text format so that text embeddings process the same room information.
@@
We assess type-constrained retrieval (nDCG@k, triplet accuracy, entity overlap), clustering quality (NMI/ARI/Silhouette), and geometric rank correlation (Spearman) against a geometry oracle. Full metric definitions and cross-type retrieval tables are in the Supplement.
Table 3: Embedding Quality Across Type-Constrained Retrieval, Clustering, and Geometric Correlation
Type-Constrained Retrieval @ $k=5$ Clustering & Geometric Correlation Method nDCG Triplet Entity Sim NMI ARI Silhouette Spearman SBM 68.2 99.9 77.0 0.726 0.437 0.141 0.184 E5-Large-v2 86.7 99.9 78.4 0.371 -0.019 0.051 0.315 BGE-Large-en-v1.5 80.1 99.9 77.0 0.358 -0.003 0.069 0.235 GTE-Large-en-v1.5 74.3 99.8 77.0 0.335 -0.002 0.067 0.133
[…]
Table˜3 (left) shows that text encoders lead on within-type ranking (nDCG: E5 86.7 vs. SBM 68.2), while SBM maintains near-perfect triplet accuracy (99.9) and comparable entity overlap (77.0). However, Table˜3 (right) reveals that SBM yields substantially better global organization (NMI 0.726 vs. 0.335–0.371, ARI 0.437 vs. $\leq$0), demonstrating nearly twice the clustering quality (Fig.˜4). This tighter type-structured embedding space—where rooms of the same category cluster more coherently—is a critical property for retrieval-augmented generation in architectural design.